# arXiv:2603.02098v1[cs.IR] 02 Mar 2026

### v1

@@ Abstract
Multimodal retrieval is the task of aggregating information from queries across heterogeneous modalities to retrieve desired targets. State-of-the-art multimodal retrieval models can understand complex queries, yet they are typically limited to two modalities: text and vision. This limitation impedes the development of universal retrieval systems capable of comprehending queries that combine more than two modalities. To advance toward this goal, we present OmniRet, the first retrieval model capable of handling complex, composed queries spanning three key modalities: text, vision, and audio. Our OmniRet model addresses two critical challenges for universal retrieval: computational efficiency and representation fidelity. First, feeding massive token sequences from modality-specific encoders to Large Language Models (LLMs) is computationally inefficient. We therefore introduce an attention-based resampling mechanism to generate compact, fixed-size representations from these sequences. This shared module is designed to maintain representational diversity and generalization capabilities while remaining sensitive to modality-specific information. Second, compressing rich omni-modal data into a single embedding vector inevitably causes information loss and discards fine-grained details. We propose Attention Sliced Wasserstein Pooling to preserve these fine-grained details, leading to improved omni-modal representations. OmniRet is trained on an aggregation of approximately 6 million query-target pairs spanning 30 datasets. We benchmark our model on 13 retrieval tasks and a MMEBv2 subset. Our model demonstrates significant improvements on composed query, audio and video retrieval tasks, while achieving on-par performance with state-of-the-art models on others. Furthermore, we curate a new Audio-Centric Multimodal Benchmark (ACM). This new benchmark introduces two critical, previously missing tasks—composed audio retrieval and audio-visual retrieval—to more comprehensively evaluate a model’s omni-modal embedding capacity. We believe our benchmark will facilitate the development of universal retrieval systems. Code and data is available at github.com/hmchuong/omniret.
@@ 1 Introduction
To address these challenges, we introduce OmniRet, the first unified framework for omni-modal embedding, capable of handling multimodal queries and targets from three primary modalities: text, vision, and audio ( Fig. 1(a)). Our framework makes two primary contributions. First, to tackle the efficiency bottleneck, we introduce an attention-based resampling module that intelligently condenses large sequences of media tokens into a compact, fixed-size set. This module is shared across modalities to enhance generalization, yet maintains modality-specific sensitivity by using separate media latents. We augment this with a novel diversity loss function, ensuring that these condensed tokens retain maximal information and thereby improve the final embedding quality. Second, to achieve high-fidelity representation without sacrificing efficiency, we propose a new pooling method called Attention Sliced Wasserstein Pooling (ASWP), inspired by PSWE [53]. By conceptualizing the set of LLM output tokens as a distribution, ASWP computes a rich embedding based on the distance to a set of learnable references. This approach preserves fine-grained, token-level information while maintaining the speed and simplicity of a single-vector system. Our extensive experiments demonstrate that OmniRet achieves superior performance on composed retrieval tasks, including those for images, videos, and audio.
@@ 3.2 Model Architecture
Attention Sliced Wasserstein Pooling. To aggregate the LLM’s output hidden states, we first condense the full sequence into a compact set of $S$ latent embeddings, $\mathbf{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{S}|\mathbf{z}\in\mathbb{R}^{D}\}$. This is done using an attention-based resampler, which is the same architecture as the Shared Media Resampler (shown in Fig. 3, left) but applied to the LLM’s output tokens.
We then pool this set $\mathbf{Z}$ using our Attention Sliced Wasserstein Pooling (ASWP) method, which is inspired by PSWE [53]. This approach, analogous to vector quantization [51, 83], is designed to preserve fine-grained information. It compares $\mathbf{Z}$ to a set of $S$ learnable reference points, $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{S}|\mathbf{x}\in\mathbb{R}^{D}\}$, across $L$ different 1D projections (projectors $\Theta=\{\theta_{1},\ldots,\theta_{L}|\theta\in\mathbb{S}^{D-1}\}$). This process computes an intermediate representation $\mathbf{Z}^{\prime}\in\mathbb{R}^{S\times L}$:
$\mathbf{Z}^{\prime}=[\psi_{1}(\mathbf{X},\mathbf{Z};\theta_{1});\ldots;\psi_{L}(\mathbf{X},\mathbf{Z};\theta_{L})]$ (7)
where $\psi_{i}(\cdot)\in\mathbb{R}^{S}$ computes the Monge coupling [69, 30] between $\mathbf{X}$ and $\mathbf{Z}$ in the 1D space defined by $\theta_{i}$.
While $\mathbf{Z}^{\prime}$ effectively preserves fine-grained details, its $S\times L>>D$ size is inefficient for a final embedding. Therefore, ASWP introduces an attention-pooling step to aggregate information from $\mathbf{Z}^{\prime}$. We use a straight-through maximum (STM) trick to create a binary attention mask $m(\mathbf{Z^{\prime}})\in\{0,1\}^{S\times L}$. For each column $\psi_{i}$ of $\mathbf{Z}^{\prime}$ (representing the $i$-th projection), we compute a soft score $\mathbf{y}=\text{softmax}(\psi_{i})\in[0,1]^{S}$. The corresponding column $m_{i}$ of the binary mask is then:
$m_{i}$ $=\text{OneHot}\left(\operatorname*{arg\,max}_{0\leq j<S}\mathbf{y}_{j}\right)-\text{StopGrad}(\mathbf{y})+\mathbf{y}$ (8)
This mask $m$ selects the single most relevant reference point for each projection. We apply it via Hadamard product: $\mathbf{V}=\mathbf{Z^{\prime}}\odot m(\mathbf{Z^{\prime}})$, where $\odot$ is the Hadamard product. The final $L$-dimensional embedding $\mathbf{h}$ is computed by a column-sum operation on $\mathbf{V}$: $\mathbf{h}_{i}=\sum_{j=1}^{S}\mathbf{V}_{ji}$. This procedure is illustrated in Fig. 3 (right). In practice, we learn the sliced references of $\mathbf{X^{\prime}}$ directly.
@@ 5.1 Implementation Details
Baselines. On the extended M-BEIR, we evaluate against dual-encoder models trained on text-binding tasks (e.g., image-to-text) and recent baselines finetuned on multimodal (image and video) data. For MMEBv2 baselines, we gather results for models with LLM backbones smaller than 4B parameters to be comparable with our model. On our benchmark with the audio modality, we choose to compare with ImageBind [16], CLAP [74], and Gemma Embedding [68]. While ImageBind [16] is a strong baseline for audio-image-video alignment, the Gemma encoding captions from QwenOmni [76] serves as an alternative solution for the composed audio retrieval task. Details of each baseline are discussed in the supplementary material.
@@ 6 Limitations and Conclusion
In conclusion, we present a novel approach to omni-modality retrieval where image, video, audio, and text are encoded into a unified embedding space by a single model. Our contributions include: (1) An innovative architecture that encodes any modality into a shared embedding space, considering both the diversity and efficiency of the extracted information; (2) A new Sliced Wasserstein embedding pooling method that overcomes the limitations of coarse-grained, single-vector embeddings; (3) A new training strategy and dataset combination that improve the model’s generalization across many retrieval tasks; (4) New ACM benchmark for evaluating interactions between audio and other modalities. Notably, OmniRet achieves superior results with 12/13 leading performance on extended M-BEIR, outstanding results on video benchmarks on MMEBv2 and promising performance on our ACM. These advancements collectively push the boundaries of omni-modality retrieval, offering more efficient and effective solutions for real-world applications like RAG or recommendation systems.
@@ 7 Additional Method Details
We leverage the instructions introduced in M-BEIR [71] and extend them to cover the specific requirements of our new datasets. We provide a complete list of all instructions used during model training in the included file, instructions.csv. Note that the first instruction in this file is the one employed during inference and testing.
@@ 8 Additional Benchmark Details
CLAP Baseline. We also benchmark against the CLAP model [74]. We adapt the fusion method from the UniIR framework [71] to handle the composed audio retrieval task. Specifically, the query is modeled as a linear combination of the query audio embedding and the modification text embedding:
@@
ImageBind Baseline. Our final baseline is the ImageBind model [16], which learns a joint embedding space across six different modalities.
  • Composed Audio Retrieval: We follow the identical experimental setup described for the CLAP baseline.

### v2

# Efficient and High-Fidelity Omni Modality Retrieval

@@ Abstract
Multimodal retrieval is the task of aggregating information from queries across heterogeneous modalities to retrieve desired targets. State-of-the-art multimodal retrieval models can understand complex queries, yet they are typically limited to two modalities: text and vision. This limitation impedes the development of universal retrieval systems capable of comprehending queries that combine more than two modalities. To advance toward this goal, we present OmniRet, the first retrieval model capable of handling complex, composed queries spanning three key modalities: text, vision, and audio. Our OmniRet model addresses two critical challenges for universal retrieval: computational efficiency and representation fidelity. First, feeding massive token sequences from modality-specific encoders to Large Language Models (LLMs) is computationally inefficient. We therefore introduce an attention-based resampling mechanism to generate compact, fixed-size representations from these sequences. This shared module is designed to maintain representational diversity and generalization capabilities while remaining sensitive to modality-specific information. Second, compressing rich omni-modal data into a single embedding vector inevitably causes information loss and discards fine-grained details. We propose Attention Sliced Wasserstein Pooling to preserve these fine-grained details, leading to improved omni-modal representations. OmniRet is trained on an aggregation of approximately 6 million query-target pairs spanning 30 datasets. We benchmark our model on 13 retrieval tasks and a MMEBv2 subset. Our model demonstrates significant improvements on composed query, audio and video retrieval tasks, while achieving on-par performance with state-of-the-art models on others. Furthermore, we curate a new Audio-Centric Multimodal Benchmark (ACM). This new benchmark introduces two critical, previously missing tasks—composed audio retrieval and audio-visual retrieval—to more comprehensively evaluate a model’s omni-modal embedding capacity. We believe our benchmark will facilitate the development of universal retrieval systems.
@@ 1 Introduction
To address these challenges, we introduce OmniRet, the first unified framework for omni-modal embedding, capable of handling multimodal queries and targets from three primary modalities: text, vision, and audio ( Fig. 1(a)). Our framework makes two primary contributions. First, to tackle the efficiency bottleneck, we introduce an attention-based resampling module that intelligently condenses large sequences of media tokens into a compact, fixed-size set. This module is shared across modalities to enhance generalization, yet maintains modality-specific sensitivity by using separate media latents. We augment this with a novel diversity loss function, ensuring that these condensed tokens retain maximal information and thereby improve the final embedding quality. Second, to achieve high-fidelity representation without sacrificing efficiency, we propose a new pooling method called Attention Sliced Wasserstein Pooling (ASWP), inspired by PSWE [52]. By conceptualizing the set of LLM output tokens as a distribution, ASWP computes a rich embedding based on the distance to a set of learnable references [53]. This approach preserves fine-grained, token-level information while maintaining the speed and simplicity of a single-vector system. Our extensive experiments demonstrate that OmniRet achieves superior performance on composed retrieval tasks, including those for images, videos, and audio.
@@ 3.2 Model Architecture
Attention Sliced Wasserstein Pooling. To aggregate the LLM output hidden states, we first compress the full sequence into a compact set of $S$ latent embeddings, $\mathbf{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{S}\mid\mathbf{z}_{i}\in\mathbb{R}^{D}\}$, using an attention-based resampler identical to the Shared Media Resampler in Fig. 3 (left), applied to the LLM tokens. Instead of average pooling $\mathbf{Z}$, which can blur fine-grained token structure, we use Attention Sliced Wasserstein Pooling (ASWP) to summarize the distribution of $\mathbf{Z}$ relative to a set of learnable reference points $\mathbf{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{S}\mid\mathbf{x}_{i}\in\mathbb{R}^{D}\}$.
Inspired by PSWE [52], ASWP compares $\mathbf{Z}$ and $\mathbf{X}$ across $L$ 1D projections $\Theta=\{\theta_{1},\ldots,\theta_{L}\mid\theta_{i}\in\mathbb{S}^{D-1}\}$. For each projection $\theta_{i}$, we project both sets onto $\theta_{i}$ and compute their 1D Monge coupling, producing
$\mathbf{Z}^{\prime}=[\psi_{1}(\mathbf{X},\mathbf{Z};\theta_{1});\ldots;\psi_{L}(\mathbf{X},\mathbf{Z};\theta_{L})]\in\mathbb{R}^{S\times L},$ (7)
where $\psi_{i}(\cdot)\in\mathbb{R}^{S}$ measures how the projected token distribution aligns with the projected references. Thus, $\mathbf{Z}^{\prime}$ can be viewed as a learnable histogram-like descriptor that preserves fine-grained distributional information.
Since $\mathbf{Z}^{\prime}$ is larger than the target embedding size, ASWP further aggregates it with a hard selection step. For each column $\psi_{i}$ of $\mathbf{Z}^{\prime}$, we compute a soft score $\mathbf{y}=\text{softmax}(\psi_{i})\in[0,1]^{S}$ and form a one-hot mask
$m_{i}^{\text{hard}}=\text{OneHot}\left(\operatorname*{arg\,max}_{0\leq j<S}\mathbf{y}_{j}\right).$ (8)
To backpropagate through this discrete choice, we use the straight-through maximum (STM) estimator:
$\tilde{m}_{i}=m_{i}^{\text{hard}}-\text{StopGrad}(\mathbf{y})+\mathbf{y}.$ (9)
In the forward pass, $\tilde{m}_{i}$ is equivalent to the binary mask $m_{i}^{\text{hard}}$, while in the backward pass it allows gradients to flow through $\mathbf{y}$. We then apply this selection by Hadamard product, $\mathbf{V}=\mathbf{Z^{\prime}}\odot\tilde{m}(\mathbf{Z^{\prime}})$, and obtain the final $L$-dimensional embedding $\mathbf{h}$ by column-wise summation: $\mathbf{h}_{i}=\sum_{j=1}^{S}\mathbf{V}_{ji}$. This procedure is illustrated in Fig. 3 (right). In practice, we learn the sliced references of $\mathbf{X^{\prime}}$ directly.
@@ 5.1 Implementation Details
Baselines. On the extended M-BEIR, we evaluate against dual-encoder models trained on text-binding tasks (e.g., image-to-text) and recent baselines finetuned on multimodal (image and video) data. For MMEBv2 baselines, we gather results for models with LLM backbones smaller than 4B parameters to be comparable with our model. On our benchmark with the audio modality, we choose to compare with ImageBind [16], CLAP [74], ViT-Lens [34], OmniBind [69] and Gemma Embedding [68]. While ImageBind, ViT-Lens, and OmniBind are strong baselines for audio-image-video alignment, the Gemma encoding captions from QwenOmni [76] serves as an alternative solution for the composed audio retrieval task. Details of each baseline are discussed in the supplementary material.
@@ 6 Limitations and Conclusion
In conclusion, we present a novel approach to omni-modality retrieval where image, video, audio, and text are encoded into a unified embedding space by a single model. Our contributions include: (1) An innovative architecture that encodes any modality into a shared embedding space, considering both the diversity and efficiency of the extracted information; (2) A new Sliced Wasserstein embedding pooling method that overcomes the limitations of coarse-grained, single-vector embeddings; (3) A new training strategy and dataset combination that improve the model’s generalization across many retrieval tasks; (4) New ACM benchmark for evaluating interactions between audio and other modalities. Notably, OmniRet achieves superior results with 12/13 leading performance on extended M-BEIR, outstanding results on video benchmarks on MMEBv2 and promising performance on ACM. These advancements collectively push the boundaries of omni-modality retrieval, offering more efficient and effective solutions for real-world applications like RAG or recommendation systems.
@@ 7 Additional Method Details
We leverage the instructions introduced in M-BEIR [71] and extend them to cover the specific requirements of our new datasets. We provide a complete list of all instructions used during model training in the code. Note that the first instruction in this file is the one employed during inference and testing.
@@ 8 Additional Benchmark Details
We conduct additional studies to assess the quality of ACM on a 300-sample subset. Since the dataset is generated by Gemini 2.5, we use GPT-4o for evaluation to avoid self-preference bias. For each sample, we score naturalness, fluency, and hallucination on a 1–5 scale. ACM achieves strong average scores of 4.4, 4.1, and 4.5, respectively. On a clean subset of 250 samples (Table 9), where all scores are at least 4, OmniRet maintains results consistent with Table 4, indicating that synthetic artifacts or distribution shifts do not favor our method.
Table 9: Performance on ACM Clean Subset.
Model A,T $\rightarrow$ A A $\rightarrow$ $V\rightarrow$ $A\rightarrow$ $I\rightarrow$ A
CLAP [12] 18.4 - - - -
ImageBind [16] 10.5 38.7 39.4 33.9 33.5
OmniRet 28.2 34.2 36.7 26.0 24.0
@@
CLAP. We also benchmark against the CLAP model [74]. We adapt the fusion method from the UniIR framework [71] to handle the composed audio retrieval task. Specifically, the query is modeled as a linear combination of the query audio embedding and the modification text embedding:
@@
ImageBind. Another baseline is the ImageBind model [16], which learns a joint embedding space across six different modalities.
  • Composed Audio Retrieval: We follow the identical experimental setup described in Eq. (12) with $\lambda=0.5$.
@@
ViT-Lens and OmniBind. Similar to ImageBind, two recent works, ViT-Lens [34] and OmniBind [69], learn a joint embedding space across multiple unimodal encoders. We follow the same task setting as ImageBind: for composed audio retrieval, we fuse audio and text embeddings, while for audio-visual retrieval, we directly use their joint embeddings.